# Sco X-1 as a continuous gravitational waves source: modelling the secular evolution using MESA

### v1

@@ Abstract
We study the prospects for detecting continuous gravitational waves from Sco X-1, and evaluate the most likely waveform- and progenitor- parameters. We achieve this by studying the evolution of different Sco X-1 progenitors and identify those that give rise to detectable signals during the accretion phase, which is what we observe now in Sco X-1. During this phase we model the spin evolution of the neutron star by two torques, the accretion torque and the gravitational wave torque. We consider two mechanisms for generating the non-axisymmetric mass distribution responsible for the gravitational wave torque due to accretion: i) magnetic mountains and ii) residual deformation after crustal breakage. Both torques are deeply intertwined with the binary evolution, which we trace from the formation of the neutron star in a binary system with a main-sequence companion. We do this with the MESA code, starting from a set of initial binary configurations.
At the LIGO-O3 sensitivity level, a magnetic-mountain ellipticity of $\varepsilon\gtrsim 10^{-6}$ is necessary for detection. The highest frequency at which we have detectable signals increases with the accretion efficiency; it can be as high as 360 Hz and presents negligible frequency variation. At 3G (Cosmic Explorer/Einstein telescope) sensitivities, less deformed Sco X-1 neutron stars, with $\varepsilon$ as small as $6\cdot 10^{-9}$, are detectable, but the waveform parameters highly depend on the binary system: the highest frequency of detectable signals spans a very broad range, between 600 Hz and 1700 Hz, strongly depending on the accretion efficiency $\eta$ and mass of the progenitor donor star $M^{d}$.
In Sco X-1-like systems with accretion efficiency $\eta\leq 30\%$, the crust of the neutron star does not break. When $\eta\in[40\%,60\%]$ only progenitors with donor mass $M^{d}\geq[1.1,1.5]$ M present crustal breakage, while if $\eta\geq 70\%$ all crusts break, independently of the progenitor parameters. In some systems the crust breaks during their Sco X-1-like phase. If Sco X-1 was one of those systems, then it would be emitting a very loud gravitational wave signal that sweeps from $\mathcal{O}(1000)$ Hz down to torque-balance frequencies of $\mathcal{O}(100)$ Hz in $150\,000\,[\varepsilon/10^{-5}]^{-2/5}$ years. We estimate the current detection probability for this type of signal to be under 1%; however, this probability increases substantially – to around 41% – with third-generation detectors.
@@ 1 Introduction
The reasoning behind Bildstein’s conjecture has to do with the fact that the spin-down torque exerted by the emission of continuous gravitational waves grows with the fifth power of the star’s angular velocity, limiting de facto the recycling frequency. Moreover, the non-axisymmetric mass distribution on the neutron star, necessary for gravitational wave emission to occur, may have its cause in the intricate process of mass accretion. Within this context, rapidly accreting systems should be the brightest gravitational wave sources, because the torque to counterbalance is correspondingly higher (details in Section 4.1). To date, Sco X-1 is the most rapidly accreting persistent low mass X-ray binary known.
Searches for continuous gravitational waves from Sco X-1 have been carried out almost continuously for about two decades. From the first search of a periodic signal in LIGO S2 data (Abbott et al., 2007) to the most recent ones in LIGO’s third observational run (Abbott et al., 2022; Whelan et al., 2023), searches have greatly improved in sensitivity and frequency coverage. Given that a detection is still missing, search results consist primarily in upper limits on the wave amplitude throughout the frequency band surveyed.
@@
This, however, does not mean accreting neutron stars do not radiate continuous gravitational waves. Several viable deformation mechanisms are indeed plausibly at work in such systems (Bildsten, 1998; Ushomirsky et al., 2000; Melatos & Payne, 2005; Morales & Horowitz, 2024, 2025). Moreover, due to its very high mass accretion rate, rather than a typical low mass X-ray binary, Sco X-1 is an outlier difficult to fit into the analysis of Andersson et al. (2005) and Patruno et al. (2012), as the authors themselves recognise. Precisely for these reasons, Sco X-1 remains one of the most interesting potential sources of continuous gravitational waves worth to study.
@@ 4.1 Torques at work on the accreting neutron star
For instance, at $M=1.65\,M_{\odot}$ and $\frac{\Omega}{2\pi}=500\,\textrm{Hz}$, the above equation gives $R_{ISCO}\approx 13\,\textrm{km}$, which is greater than the radius at any mass for the AP3 equation of state.
We therefore define
@@ 5.2 Crust failure
$\frac{\Omega_{K}}{2\pi}=1716\sqrt{\frac{M/1.4M_{\odot}}{\left(R/10\textrm{km}\right)^{3}}}-189\,\textrm{Hz}.$ (26)
The deformation left by crust breakage is not predictable. We consider a grid of 19 ellipticities log-uniformly distributed in the range $\log\varepsilon\in[-9,-4.5]$
@@ 6.1 Sco X-1 progenitors
We do not focus on this aspect of our simulations, as it has been well discussed in Van & Ivanova (2021), and our results are largely consistent with theirs. The major difference is that we find Sco X-1 progenitors for all values of accretion efficiency $\eta$ other than for $\eta=20\%$, which is too low to result in compatible mass accretion rates. Conversely, Van & Ivanova fix the accretion efficiency exactly at $20\%$, but allow for Sco X-1 to have accretion rates as low as $0.32\cdot 10^{-8}$, significantly below our lower limit of Equation 3.
@@ 6.2.1 Effects of accretion efficiency and mass of the donor star
Figure 2: The three rows of plots differ for the value of the accretion efficiency: 30%, 50% and 80%, and refer to the same progenitor system defined by: $P=1.91\,\textrm{d},M^{d}=1.25\,M_{\odot}$. On the left we have the spin frequency evolution of the neutron star in Sco X-1 in the case of high deformation (solid blue curve) and small deformation (dashed blue curve), in both cases due to magnetic confinement of accreted matter. The twin $y$ axis shows the moment of inertia as a function of time (orange curve). We highlight the evolutionary stage where the system resembles Sco X-1 with scatter markers superimposed on the curves (in this example, in green). On the right we plot the continuous gravitational waves emitted by the highly and weakly deformed neutron star (same as the left-side plots). We compare these with the most recent gravitational wave amplitude upper limits $h_{0}^{ul}$ (Abbott et al. (2022), solid grey curve), as well as with the nominal torque-balance condition discussed in Section 4.1. We remind that the nominal torque balance condition assumes $M=1.4\,M_{\odot}$, $R_{i}=10\,\textrm{km}$ and $\dot{M}_{\textrm{a}}=3.1\cdot 10^{-8}M_{\odot}\cdot\textrm{yr}^{-1}$.
@@ 6.2.2 Constraints on magnetic mountains
For magnetic mountains, the magnitude of the ellipticity depends on the magnetic flux ratio $b$, which can hence be constrained based on null-detection results. Assuming Sco X-1 to be in nominal torque balance (Equation 23), the gravitational wave frequencies $f$ at which a search could have detected the signal are
@@
where $h_{0}^{tb}$ is the torque balance amplitude and $h_{0}^{ul}$ is the observational upper limit value. The $h_{0}^{ul}$ can be translated in a lower limit on $b$, using Equation 25 with $\Delta M\geq 8M_{c}/b^{2}$. We consider the most recent gravitational wave amplitude upper limit results from Abbott et al. (2022), and find that the detectability region ($h_{0}^{tb}\geq h_{0}^{ul}$) falls roughly between 40 and 200 Hz. In this region, the null detection results yield $b\gtrsim 18$, corresponding to $\varepsilon\leq 3.9\cdot 10^{-6}$.
The estimate above equates the detectability level to nominal torque balance. We can also determine the detectability directly based on our results: a signal stemming from a simulated system is detectable if its frequency and frequency time derivative fall within the parameter space covered by Abbott et al. (2022) and if its amplitude is greater than amplitude upper limits at the signal’s frequency.
The frequency variability of Sco X-1 can be both short-term – with timescales between months to thousands of years – and long-term – with timescales between hundreds of thousands to millions of years. Our study is not suited to forecast short-term variability on Sco X-1. Firstly, because we do not model very short-term effects such as spin-wandering (Mukherjee et al., 2018) characterised by timescales of the order of days to months. Secondly, the mass transfer rates of systems simulated with MESA present numerical noise affecting the spin frequency time derivative on timescales $\lesssim 10^{5}\,\textrm{yr}$, therefore impacting the higher end of timescales pertaining to short-term variability.
@@
The $\dot{f}$ range covered by Abbott et al. (2022), is not given in the paper. We estimate the latter by using Equation 86 of Leaci & Prix (2015), with $t_{s}=T_{\textrm{obs}}$, consistent with having assumed that the timescale during which the signal varies is equal to the total observing time $T_{\textrm{obs}}$:
@@
where $\Delta f$ is the frequency resolution. From Table 3 of Abbott et al. (2022) the frequency resolution $\Delta f\geq 5.3\times 10^{-5}\,\textrm{Hz}$, therefore using $T_{\textrm{obs}}=3\cdot 10^{7}\,\textrm{s}$ which is the duration of the LIGO O3 run, we find $\mid\dot{f}_{\textrm{max}}\mid\geq 2.4\cdot 10^{-12}\,\textrm{Hz}\cdot\textrm{s}^{-1}$.
All the simulated signals emitting gravitational waves at the current detectable level have $6\cdot 10^{-13}\leq\dot{f}^{\textrm{lt}}/[\textrm{Hz}\cdot\textrm{s}^{-1}]\leq 2.8\cdot 10^{-13}$ and frequencies much below 1600 Hz, therefore they all lie within the parameter space coverage of Abbott et al. (2022).
@@
The primary reason why we find detectable signals at frequencies significantly higher than the $\approx 200\,\textrm{Hz}$ estimated by means of nominal torque balance is due to the value of the inner radius $R_{i}$ in Eq. 23, assumed to be $10\,\textrm{km}$ in the nominal torque balance condition, but that we find to be typically between $17\leq R_{i}/[\textrm{km}]\leq 30$ in our simulations.
@@ 6.2.3 Third generation detectors
Let us consider third-generation ground-based detectors. Following the procedure already adopted and amply discussed in Pagliaro et al. (2023) and Pagliaro et al. (2025b), we imagine to replicate the search done by Abbott et al. (2022), but on more sensitive data. This search reaches a sensitivity depth (Behnke et al., 2015) of $\mathcal{D}\approx 62\,\textrm{Hz}^{-1/2}$ using both LIGO detectors. The amplitude of the weakest detectable signal at frequency $f$ is therefore [^2] See Section 2.1 of Pagliaro et al. (2025b) for a detailed discussion and derivation.
@@
All the simulated signals emitting gravitational waves at $3G$ detectable level have $6.1\cdot 10^{-13}\leq\dot{f}^{\textrm{lt}}\leq 2.2\cdot 10^{-12}$. We can conclude that, as far as magnetic mountains are concerned, the long-term spin frequency evolution does not impact the sensitivity of gravitational wave searches like Abbott et al. (2022), at least for observation times of $\approx 1\,\textrm{yr}$.
@@ 6.3 Crustal breakage
We next present results for the case where the crusts fails to sustain centrifugal stresses and brakes, leaving behind a residual deformation $\varepsilon$ (Section 5.2). Gravitational wave emission only begins after the breakage. This is illustrated in Figure 5, where we can observe that the large assumed ellipticity $\varepsilon=10^{-5}$ is responsible for a dramatic change in the spin frequency evolution after the breakage. In our simulation we consider a range of $\varepsilon\in[10^{-9},3.16\times 10^{-5}]$.
@@
The breakage, and hence the gravitational wave emission, is predicated upon the neutron star reaching the angular velocity $\Omega^{\textrm{br}}=0.45\cdot\Omega^{K}$. This introduces a strong dependency on the accretion efficiency because a larger accretion efficiency produces a more efficient spin-up, as discussed in Section 6.2.1. We find that for $\eta\leq 30\%$ no binary configuration compatible with Sco X-1 manages to spin the neutron star up to $\Omega^{\textrm{br}}$. For $\eta=40\%$ crustal failure only happens for Sco X-1 systems with donor mass exceeding $1.52\,M_{\odot}$ (see Figure 6 ). As the accretion efficiency increases more and more Sco X-1 systems present crustal breakage, as lower mass donors are able to sufficiently spin up the star. For $\eta\geq 70\%$, all progenitors lead to crustal breakage (see Table 2).
Figure 5: Spin evolution (solid lines, left plot) and concurrent gravitational wave emission (right plot) of a Sco X-1 neutron star going through crustal failure, which leaves behind a residual ellipticity of $\varepsilon=10^{-5}$, for three values of the accretion efficiency: 30%, 40% and 50%; among these, only $\eta=50\%$ leads to crustal failure therefore to continuous gravitational wave emission. The dashed lines on the left plot represent $45\%$ of the Keplerian break-up spin frequency as a function of time. The gravitational wave "trace" of the right plot is color-coded with the instantaneous gravitational wave frequency time derivative. It takes $\approx 250\,\textrm{yr}$ to spin-down from A, the point of crustal breakage, to B, where spin frequency is half of the break-up frequency. It takes further $\approx 3500\,\textrm{yr}$ to reach C, where the spin frequency is reduced by another factor of 2. In D the neutron star is in torque balance, persisting there for slightly less than a million years before leaving the Sco X-1 compatibility state. The progenitor system considered here is the same as the one considered in Figure 2.
@@
Table 2: Donor masses of Sco X-1 progenitor systems leading to crustal breakage. In brackets, the subset of these for which the neutron star’s crust breaks during the Sco X-1 compatible phase.
@@
One of the distinct features of this deformation mechanism is the strength of emitted gravitational waves, which reaches amplitudes almost two orders of magnitude higher than the magnetic mountain case, as shown in Figure 8, which serves as a summary figure for this deformation case. This is due to the fact that the gravitational wave emission starts at $f^{\textrm{br}}=\Omega^{\textrm{br}}/\pi$ which is high (above $\approx 1000$ Hz) and the gravitational wave amplitude is quadratic in ff(Equation 22). If the system is in the Sco X-1 compatible phase when the crust breaks, this means that the entire signal could be observed, as it sweeps from $f^{\textrm{br}}$ to the much lower torque balance frequency. It should however be noted that the high-frequency/loud gravitational wave phase is very short (tens of thousands of years) compared to the low-frequency/weak gravitational wave phase (hundreds of thousands to millions of years), making it unlikely to observe such signal during its loudest phases. The rapid spindown in the loud phase also falls outside the range covered by the most sensitive searches like Abbott et al. (2022). It is nonetheless covered by the radiometer searches, whose sensitivity in Abbott et al. (2021) would be sufficient to detect all signals in the high-frequency/loud gravitational wave phase if the ellipticity after crustal breakage is $\varepsilon\geq 10^{-6}$, highlighting the importance of robust detection techniques.

### v2

@@ Abstract
[…]
@@ 1 Introduction
The reasoning behind Bildsten’s conjecture has to do with the fact that the spin-down torque exerted by the emission of continuous gravitational waves grows with the fifth power of the star’s angular velocity, limiting de facto the recycling frequency. Moreover, the non-axisymmetric mass distribution on the neutron star, necessary for gravitational wave emission to occur, may have its cause in the intricate process of mass accretion. Within this context, rapidly accreting systems should be the brightest gravitational wave sources, because the torque to counterbalance is correspondingly higher (details in Section 4.1). To date, Sco X-1 is the most rapidly accreting persistent low mass X-ray binary known.
Searches for continuous gravitational waves from Sco X-1 have been carried out almost continuously for about two decades. From the first search of a periodic signal in LIGO S2 data (Abbott et al., 2007) to the most recent ones in LIGO’s third observational run (Abbott et al., 2022b, a; Whelan et al., 2023; Vargas & Melatos, 2025), searches have greatly improved in sensitivity and frequency coverage. Given that a detection is still missing, search results consist primarily in upper limits on the wave amplitude throughout the frequency band surveyed.
@@
This, however, does not mean accreting neutron stars do not radiate continuous gravitational waves. Several viable deformation mechanisms are indeed plausibly at work in such systems (Bildsten, 1998; Andersson et al., 1999; Ushomirsky et al., 2000; Melatos & Payne, 2005; Morales & Horowitz, 2024, 2025). Moreover, due to its very high mass accretion rate, rather than a typical low mass X-ray binary, Sco X-1 is an outlier difficult to fit into the analysis of Patruno et al. (2012) for instance, as the authors themselves recognise. Precisely for these reasons, Sco X-1 remains one of the most interesting potential sources of continuous gravitational waves worth to study.
@@ 4.1 Torques at work on the accreting neutron star
For instance, at $M=1.65\,M_{\odot}$ and $\frac{\Omega}{2\pi}=500\,\textrm{Hz}$, the above equation gives $R_{ISCO}\approx 13\,\textrm{km}$, which is greater than the radius at any mass for the AP3 equation of state. We therefore define
@@ 5.2 Crust failure
$\frac{\Omega_{K}}{2\pi}=1716\sqrt{\frac{M/1.4M_{\odot}}{\left(R/10\textrm{km}\right)^{3}}}-189\ [\textrm{Hz}].$ (26)
The deformation left by crust breakage is not predictable. We consider a grid of 19 ellipticities log-uniformly distributed in the range $\log\varepsilon\in[-9,-4.5]$. The lower bound of $10^{-9}$ is chosen according to a possible indication of a minimum ellipticity among recycled pulsars (Woan et al., 2018); while the upper bound of $\varepsilon\approx 3.16\times 10^{-5}$ is chosen so that the newly formed mountain is still small enough to be long-lived (Horowitz & Kadau, 2009; Morales & Horowitz, 2022).
@@ 6.1 Sco X-1 progenitors
Our first task is to identify the set of Sco X-1 progenitor systems. We do not focus on this aspect of our simulations, as it has been well discussed in Van & Ivanova (2021), and our results are largely consistent with theirs. The major difference is that we find Sco X-1 progenitors for all values of accretion efficiency $\eta$ other than for $\eta=20\%$, which is too low to result in compatible mass accretion rates. Conversely, Van & Ivanova fix the accretion efficiency exactly at $20\%$, but allow for Sco X-1 to have accretion rates as low as $0.32\cdot 10^{-8}$, significantly below our lower limit of Equation 3.
@@ 6.2.1 Effects of accretion efficiency and mass of the donor star
Figure 2: The three rows of plots differ for the value of the accretion efficiency: 30%, 50% and 80%, and refer to the same progenitor system defined by: $P=1.91\,\textrm{d},M^{d}=1.25\,M_{\odot}$. On the left we have the spin frequency evolution of the neutron star in Sco X-1 in the case of high deformation (solid blue curve) and small deformation (dashed blue curve), in both cases due to magnetic confinement of accreted matter. The twin $y$ axis shows the moment of inertia as a function of time (orange curve). We highlight the evolutionary stage where the system resembles Sco X-1 with scatter markers superimposed on the curves (in this example, in green). On the right, we plot the gravitational wave amplitude $h_{0}(f)$ emitted by the highly and weakly deformed neutron star (same as the left-side plots). We compare it with the most recent gravitational wave amplitude upper limits $h_{0}^{ul}$ (Abbott et al. (2022b), solid grey curve), as well as with the nominal torque-balance condition introduced in Section 4.1. The $h_{0}(f)$ curves present a "turning point" in the case of high ellipticity: very early on, the system reaches torque balance, so the frequency stops increasing. However, since the mass accretion rate $\dot{M}_{\textrm{a}}$ decreases over time, the accretion torque also decreases, and the neutron star begins to slowly spin down. Even though the moment of inertia II increases, since $h_{0}\propto I\,f^{2}$, overall $h_{0}$ decreases, creating the “turning point".
@@ 6.2.2 Constraints on magnetic mountains
For magnetic mountains, the magnitude of the ellipticity depends on the magnetic flux ratio $b$, which can hence be constrained based on null-detection results. We first review the most common way to assess detectability in the absence of a population study, and then present our results.
Standard detectability: the most frequently used way to assess detectability of a signal at a certain frequency by a given search, is to assume that Sco X-1 is in torque balance (Equation 23), so that the gravitational wave frequencies $f$ at which the signal is detectable are
@@
where $h_{0}^{tb}$ is the torque balance amplitude and $h_{0}^{ul}$ is the observational upper limit value. The $h_{0}^{ul}$ can then be translated in a lower limit on $b$, using Equation 25 with $\Delta M\geq 8M_{c}/b^{2}$.
One often finds the torque balance amplitude computed from Equation 23 by assuming nominal values for $M,R_{i}$ and $\dot{M}_{a}$, most often taken equal to the reference scaling ones used in our equation. With this convention and considering the most recent gravitational wave amplitude upper limit results from Abbott et al. (2022b), we find that the detectability region ($h_{0}^{tb}\geq h_{0}^{ul}$) falls roughly between 40 and 200 Hz. In this region, the null detection results yield $b\gtrsim 18$, corresponding to $\varepsilon\leq 3.9\cdot 10^{-6}$.
Direct determination: based on our results we can determine the detectability directly: a signal stemming from a simulated system is detectable if its frequency and frequency time derivative fall within the parameter space covered by Abbott et al. (2022b) and if its amplitude is greater than amplitude upper limits at the signal’s frequency.
The frequency variability of Sco X-1 can be both short-term – with timescales ranging from months to thousands of years – and long-term – with timescales ranging from hundreds of thousands to millions of years. Our study is not suited to forecast short-term variability on Sco X-1. Firstly, because we do not model very short-term effects such as spin-wandering (Mukherjee et al., 2018) characterised by timescales of the order of days to months. Secondly, the mass transfer rates of systems simulated with MESA present numerical noise affecting the spin frequency time derivative on timescales $\lesssim 10^{5}\,\textrm{yr}$, therefore impacting the higher end of timescales pertaining to short-term variability.
@@
The $\dot{f}$ range covered by Abbott et al. (2022b), is not given in the paper. We estimate the latter by using Equation 86 of Leaci & Prix (2015), with $t_{s}=T_{\textrm{obs}}$, consistent with having assumed that the timescale during which the signal varies is equal to the total observing time $T_{\textrm{obs}}$:
@@
where $\Delta f$ is the frequency resolution. From Table 3 of Abbott et al. (2022b) the frequency resolution $\Delta f\geq 5.3\times 10^{-5}\,\textrm{Hz}$, therefore using $T_{\textrm{obs}}=3\cdot 10^{7}\,\textrm{s}$ which is the duration of the LIGO O3 run, we find $\mid\dot{f}_{\textrm{max}}\mid\geq 2.4\cdot 10^{-12}\,\textrm{Hz}\cdot\textrm{s}^{-1}$.
All the simulated signals emitting gravitational waves at the current detectable level have $6\cdot 10^{-13}\leq\dot{f}^{\textrm{lt}}/[\textrm{Hz}\cdot\textrm{s}^{-1}]\leq 2.8\cdot 10^{-13}$ and frequencies much below 1600 Hz, therefore they all lie within the parameter space coverage of Abbott et al. (2022b).
@@
Detectability established by direct comparison of the upper limits with our simulated Sco X-1 population yields detectable signals at frequencies higher than $\approx$ 200 Hz, which is instead the highest currently detectable frequency based on nominal torque balance. The primary reason is that in nominal torque balance $R_{i}=10$ km, whereas in our simulations we find $17\leq R_{i}/[\textrm{km}]\leq 30$. As explained at the beginning of this Section, the nominal torque balance condition is often used as a standard benchmark to gauge to what extent null search results are astrophysically relevant. Our numerical results confirm what is already understood in searches, namely that the gravitational wave amplitude calculated assuming the nominal torque balance condition is a conservative estimation of the strength of the gravitational waves emitted by Sco X-1, assuming this is in torque balance.
@@ 6.2.3 Third generation detectors
Let us consider third-generation ground-based detectors. Following the procedure already adopted and amply discussed in Pagliaro et al. (2023) and Pagliaro et al. (2025b), we imagine to replicate the search done by Abbott et al. (2022b), but on more sensitive data. This search reaches a sensitivity depth (Behnke et al., 2015) of $\mathcal{D}\approx 62\,\textrm{Hz}^{-1/2}$ using both LIGO detectors. The amplitude of the weakest detectable signal at frequency $f$ is therefore [^2] See Section 2.1 of Pagliaro et al. (2025b) for a detailed discussion and derivation.
@@
All the simulated signals emitting gravitational waves at $3G$ detectable level have $6.1\cdot 10^{-13}\leq\dot{f}^{\textrm{lt}}\leq 2.2\cdot 10^{-12}$. We can conclude that, as far as magnetic mountains are concerned, the long-term spin frequency evolution does not impact the sensitivity of gravitational wave searches like Abbott et al. (2022b), at least for observation times of $\approx 1\,\textrm{yr}$.
@@ 6.3 Crustal breakage
We next present results for the case where the crust fails to sustain centrifugal stresses and breaks, leaving behind a residual deformation $\varepsilon$ (Section 5.2). Gravitational wave emission only begins after the breakage. This is illustrated in Figure 5, where we can observe that the large assumed ellipticity $\varepsilon=10^{-5}$ is responsible for a dramatic change in the spin frequency evolution after the breakage. In our simulation we consider a range of $\varepsilon\in[10^{-9},3.16\times 10^{-5}]$.
@@
The breakage, and hence the gravitational wave emission, is predicated upon the neutron star reaching the angular velocity $\Omega^{\textrm{br}}=0.45\cdot\Omega^{K}$. This introduces a strong dependency on the accretion efficiency because a larger accretion efficiency produces a more efficient spin-up, as discussed in Section 6.2.1. We find that for $\eta\leq 30\%$ no binary configuration compatible with Sco X-1 manages to spin the neutron star up to $\Omega^{\textrm{br}}$. For $\eta=40\%$ crustal failure only happens for Sco X-1 systems with donor mass exceeding $1.55\,M_{\odot}$ (see Figures 6 and 8). As the accretion efficiency increases, more and more Sco X-1 systems present crustal breakage, as lower mass donors are able to sufficiently spin up the star. For $\eta\geq 70\%$, all progenitors lead to crustal breakage (see Table 2).
Figure 5: Spin evolution (solid lines, left plot) and concurrent gravitational wave emission (right plot) of a Sco X-1 neutron star going through crustal failure, which leaves behind a residual ellipticity of $\varepsilon=10^{-5}$, for three values of the accretion efficiency: 30%, 40% and 50%; among these, only $\eta=50\%$ leads to crustal failure therefore to continuous gravitational wave emission. The dashed lines on the left plot represent $45\%$ of the Keplerian break-up spin frequency as a function of time. The gravitational wave "trace" of the right plot is color-coded with the instantaneous gravitational wave frequency time derivative. It takes $\approx 250\,\textrm{yr}$ to spin-down from A, the point of crustal breakage, to B, where spin frequency is half of the break-up frequency. It takes further $\approx 3500\,\textrm{yr}$ to reach C, where the spin frequency is reduced by another factor of 2. In D, reached from C in $\approx 200\,000\,\textrm{yr}$, the neutron star is in torque balance, persisting there for slightly less than a million years before leaving the Sco X-1 compatibility state. The progenitor system considered here is the same as the one considered in Figure 2.
@@
Table 2: Donor masses of Sco X-1 progenitor systems leading to crustal breakage. In brackets, the subset of these for which the neutron star’s crust breaks during the Sco X-1 compatible phase, therefore leading to the high-frequency/loud emission of gravitational waves as discussed in Section 6.3.
@@
One of the distinct features of this deformation mechanism is the strength of emitted gravitational waves, which reaches amplitudes almost two orders of magnitude higher than the magnetic mountain case, as shown in Figure 8, which serves as a summary figure for this deformation case. This is due to the fact that the gravitational wave emission starts at $f^{\textrm{br}}=\Omega^{\textrm{br}}/\pi$ which is high (above $\approx 1000$ Hz) and the gravitational wave amplitude is quadratic in $f$ (Equation 22). If the system is in the Sco X-1 compatible phase when the crust breaks, this means that the entire signal could be observed, as it sweeps from $f^{\textrm{br}}$ to the much lower torque balance frequency. It should however be noted that the high-frequency/loud gravitational wave phase is very short (tens of thousands of years) compared to the low-frequency/weak gravitational wave phase (hundreds of thousands to millions of years), making it unlikely to observe such signal during its loudest phases. The rapid spindown in the loud phase also falls outside the range covered by the most sensitive searches like Abbott et al. (2022b). It is nonetheless covered by the radiometer searches, whose sensitivity in Abbott et al. (2021) would be sufficient to detect all signals in the high-frequency/loud gravitational wave phase if the ellipticity after crustal breakage is $\varepsilon\geq 10^{-6}$, highlighting the importance of robust detection techniques.